OpenStudy (anonymous):

The function f is defined by f(x) ≡ 2 + ln (3x − 2), x ∈ , x > 2 3 . (a) Find the exact value of ff(1). (2) (b) Find an expression for f −1 (x).

OpenStudy (anonymous):

part a) f(1)= 2+ln1 so ff(1)= 2+ln(3(2+ln1)-2) = 2+ln 4

geerky42 (geerky42):

I don't think so because f(x) is defined 2 + ln (3x − 2) for x>2, so f(1) could be different. Not really sure, though

OpenStudy (anonymous):

what about part b? I think im right for part a I just cant do part b :/

OpenStudy (anonymous):

oh wait.... y=2+ln(3x-2) so y-2=ln(3x-2) e^(y-2)=3x-2 1/3(e^(y-2)+2)=x so inverse function = 1/3(e^(x-2)+2)